# Existence of solution to a system of PDEs modeling the crystal growth inside lithium batteries

### v2

@@ 3.9. Lemma (energy estimates for the concentration)
$\frac{\tau}{2}\|d_{\tau}c^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\frac{1}{2}d_{\tau}\|c^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\!\!\!$ $+$ $\!\!\!D_{\min}|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{1}{2}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\frac{1}{2}\|c^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}$ (3.51)
$+C|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}+\frac{D_{\min}}{2}|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}.$
After rearranging (3.51), multiplying the resulting inequality by $\tau$ and summing over $k=1,2,...,l$, for any $l\leq K$, we have
@@
12($\frac{1}{2}\left(1-\tau\right)\|c^{l}\|^{2}_{\operatorname{L}_{2}(\Omega)}\!\!\!$ $+$ $\!\!\!\tau\frac{D_{\min}}{2}\sum_{k=1}^{l}|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leq C\Big{(}1+\tau\sum_{k=1}^{l}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}$ (3.52)
$+\leavevmode\nobreak\ \tau\sum_{k=1}^{l}|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}+\tau\sum_{k=1}^{l-1}\|c^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}\Big{)}.$
@@
$\begin{split}\frac{1}{4}\max_{l=1,...,K}\|c^{l}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau\frac{D_{\min}}{2}\ \sum_{k=1}^{K}|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \\ \leq C\Big{(}1+\tau\sum_{k=1}^{K}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau\sum_{k=1}^{K}|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\Big{)}\operatorname{e}\!^{\frac{T}{2}}\end{split}$ (3.53)
and from lemmas 3.5 and 3.8 we conclude
@@
$(d_{\tau}c^{k},\chi)_{((\operatorname{H}^{1}(\Omega))^{\prime},\operatorname{H}^{1}(\Omega))}\leq\|D(u^{k})\nabla c^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\nabla\chi\|_{\operatorname{L}_{2}(\Omega)}+\|d_{\tau}u^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\chi\|_{\operatorname{L}_{2}(\Omega)}\\ +\|D_{1}(u^{k},c^{k-1})\nabla\phi^{k}\|_{\operatorname{L}_{2}(\Omega)}|\nabla\chi\|_{\operatorname{L}_{2}(\Omega)}\\ \leq C(|c^{k}|_{\operatorname{H}^{1}(\Omega)}+|\phi^{k}|_{\operatorname{H}^{1}(\Omega)}+\|d_{\tau}u^{k}\|_{\operatorname{L}_{2}(\Omega)})\|\chi\|_{\operatorname{H}^{1}(\Omega)}$ (3.55)
@@
$\|d_{\tau}c^{k}\|_{(\operatorname{H}^{1}(\Omega))^{\prime}}^{2}\leq C(|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}+|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}+\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}).$ (3.56)
The required result follows after multiplying by $\tau$, summing over $k=1,2,...,K$ and noting (3.54) and lemmas 3.5 and 3.8. ∎
Proof  of Theorem 3.1 Theorem 3.1 is a direct consequence of lemmas 3.4, 3.6 and 3.7.   ∎
@@ 4. Weak convergence of the limits
$u^{-}_{\tau}(t)=u^{k-1},\quad u^{+}_{\tau}(t)=u^{k}\text{ for }t_{k-1}<t\leq t_{k},$ (4.1)
and do the same $c^{\pm}_{\tau}(t),\hat{c}_{\tau}(t),\phi^{\pm}_{\tau}(t),\hat{\phi}_{\tau}(t)$. Also use $u_{\tau}$ to denote the interpolate
@@
Next recalling (3.2), the terms on the left hand side of (3.1) can be reformulated as
$\begin{split}((m(c^{k-1}))^{-}h_{1}^{\prime}(u^{k}),\chi)+((m(c^{k-1}))^{+}h_{2}^{\prime}(u^{k}),\chi)=30\left(m(c^{k-1})u^{k-1}u^{k}(u^{k}-1)^{2},\chi\right)\\ +30\left((m(c^{k-1}))^{+}(u^{k}-1)u^{k}(u^{k-1}-u^{k}),\chi\right).\end{split}$ (4.3)
@@
$(\sigma(u^{+}_{\tau}(t))\nabla\phi^{+}_{\tau}(t),\nabla\eta)=-(\hat{u}_{\tau}^{\prime}(t)+\sigma^{\prime}(u_{\tau}^{+}(t))\partial_{x}u_{\tau}^{+}(t),\eta),$ (4.4c)
@@
$\begin{split}({\mathcal{F}}(c_{\tau}^{-}(t),u_{\tau}^{-}(t),u_{\tau}^{+}(t)),\chi):=30\left(m(c_{\tau}^{-})u_{\tau}^{-}(t)u_{\tau}^{+}(t)(u_{\tau}^{+}(t)-1)^{2},\chi\right)\\ +30\left((m(c_{\tau}^{-}(t)))^{+}(u_{\tau}^{+}(t)-1)u_{\tau}^{+}(t)(u_{\tau}^{-}(t)-u_{\tau}^{+}(t)),\chi\right).\end{split}$
@@ 4.1. Lemma (stability bounds)
The first and second bounds in (4.5) follow from lemmas 3.2 and 3.5. For the last bound in (4.5), we note from (4.2) that $\partial_{t}\hat{u}_{\tau}(t)=d_{\tau}u^{k}$ on $(t_{k-1},t_{k})$, $k=1,2,...,K$ and thus the result follows from Lemma 3.5.
The bounds in (4.6) and (4.7) follow using similar arguments and the results of lemmas 3.9 and 3.8 respectively. ∎
@@ 4.2. Lemma (weak compactness)
There exist $(u,c,\phi)$ and $(\partial_{t}u,\partial_{t}c)$ such that for a sequence $(\tau_{n})_{n\in\mathbb{N}}$ of positive numbers with $\tau_{n}\to 0$ as $n\to\infty$, we have the following for $1\leq p<\infty$ if dim $\text{dim }\Omega=2$ and $61\leq p<$ 6 if dim $\text{dim }\Omega=3$, up to a subsequence, which for simplicity of presentation we again denote by $\tau_{n}$,
@@ 4.3. Lemma (strong convergence)
For $u\in\operatorname{L}_{\infty}([0,T];\operatorname{L}_{\infty}(\Omega))$ and $c\in\operatorname{L}_{2}([0,T];\operatorname{L}_{q}(\Omega))$, with $1\leq q<\infty$ if dim $\text{dim }\Omega=2$ and $31\leq q<$ 3 if dim $\text{dim }\Omega=3$, we have the following, for $1\leq p<\infty$ if dim $\text{dim }\Omega=2$ and $61\leq p<$ 6 if dim $\text{dim }\Omega=3$, as $\tau_{n}\to 0$
@@
$D_{1}(u^{+}_{\tau_{n}},c^{-}_{\tau_{n}})\to D_{1}(u,c)\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{q}(\Omega)),$ (4.9f)
$30m(c_{\tau}^{-})u_{\tau}^{-}u_{\tau}^{+}(u_{\tau}^{+}-1)^{2}\to m(c)h^{\prime}(u)\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{q}(\Omega)),$ (4.9g)
@@
Noting that $g^{\prime}(u),D(u),\sigma(u)$ are polynomials in $u$ and $m(c)$ is linear in $c$, the proofs of (4.9a) – (4.9e) follow from Lemma 4.1, (4.8a), (4.8d), (4.8h) and (4.8h). By the triangle inequality, the generalized Hölder inequality with $1/q=1/p+1/p^{\prime}$, (4.6) and (3.9) we obtain
$\|D(u_{\tau}^{+})c_{\tau}^{-}-D(u)c\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{q}(\Omega))}\\ =\|(D(u_{\tau}^{+})-D(u))c_{\tau}^{-}+D(u)(c_{\tau}^{-}-c)\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{q}(\Omega))}\\ \leq\|c_{\tau}^{-}\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p^{\prime}}(\Omega))}\|D(u_{\tau}^{+})-D(u)\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega))}\\ +\|D(u)\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p^{\prime}}(\Omega))}\|c_{\tau}^{-}-c\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega))}\\ \leq C(\|D(u_{\tau}^{+})-D(u)\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega))}+\|c_{\tau}^{-}-c\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega))})$ (4.10)
and (4.9f) then follows by noting the definition of $D_{1}(u,c)$ in (2.10), and using (4.9c) and (4.8i). Adopting similar techniques and noting from (2.12) that $h^{\prime}(u)=30u^{2}(u-1)^{2}$, from (3.11), (4.5), (4.9b) and (4.9e) we conclude (4.9g).
@@ 4.4. Lemma (weak convergence)
For $D(u)\nabla c$, $D_{1}(u,c)\nabla\phi$, $\sigma(u)\nabla\phi$, $\boldsymbol{{A}}(\nabla u)\nabla u\in\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d})$ we have the following as $\tau_{n}\to 0$
$D(u^{+}_{\tau_{n}})\nabla c^{+}_{\tau_{n}}\rightharpoonup D(u)\nabla c\text{ in }\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d}),$ (4.12a)
$D_{1}(u^{+}_{\tau_{n}},c^{-}_{\tau_{n}})\nabla\phi^{+}_{\tau_{n}}\rightharpoonup D_{1}(u,c)\nabla\phi\text{ in }\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d}),$ (4.12b)
$\sigma(u^{+}_{\tau_{n}})\nabla\phi_{\tau_{n}}^{+}\rightharpoonup\sigma(u)\nabla\phi\text{ in }\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d}),$ (4.12c)
@@
From Lemma 4.1 we have
$\int_{0}^{T}\int_{\Omega}D(u^{+}_{\tau_{n}})\nabla c^{+}_{\tau_{n}}\cdot\nabla c^{+}_{\tau_{n}}=\int_{0}^{T}\int_{\Omega}|D^{1/2}(u^{+}_{\tau_{n}})\nabla c^{+}_{\tau_{n}}|^{2}\\ \leq\|D^{1/2}(u^{+}_{\tau_{n}})\nabla c^{+}_{\tau_{n}}\|_{\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d})}^{2}\leq D_{\max}\|c^{+}_{\tau_{n}}\|_{\operatorname{L}_{2}([0,T];\operatorname{H}^{1}(\Omega))}^{2}\leq C$ (4.13)
and hence it follows that there exists a $\xi_{1}\in\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d})$ such that
$D(u^{+}_{\tau_{n}})\nabla c^{+}_{\tau_{n}}\rightharpoonup\xi_{1}\text{ in }\operatorname{L}_{2}([0,T];(\operatorname{L}_{2}(\Omega))^{d}).$ (4.14)
We now show that $\xi_{1}=D(u)\nabla c$, to this end we note that
$\Big{|}\int_{0}^{T}\int_{\Omega}\Big{(}D(u_{\tau_{n}}^{+})\nabla c_{\tau_{n}}^{+}-D(u)\nabla c\Big{)}\cdot\nabla\psi\Big{|}\\ =\Big{|}\int_{0}^{T}\int_{\Omega}\Big{(}D(u_{\tau_{n}}^{+})\nabla c_{\tau_{n}}^{+}-D(u)\nabla c_{\tau_{n}}^{+}+D(u)\nabla c_{\tau_{n}}^{+}-D(u)\nabla c\Big{)}\cdot\nabla\psi\Big{|}\\ \leq\Big{|}\int_{0}^{T}\int_{\Omega}D(u)\Big{(}\nabla c_{\tau_{n}}^{+}-\nabla c\Big{)}\cdot\nabla\psi\Big{|}+\Big{|}\int_{0}^{T}\int_{\Omega}\Big{(}D(u^{+}_{\tau_{n}})-D(u)\Big{)}\nabla c_{\tau_{n}}^{+}\cdot\nabla\psi\Big{|}\\ \leq\Big{|}\int_{0}^{T}\int_{\Omega}D(u)\Big{(}\nabla c_{\tau_{n}}^{+}-\nabla c\Big{)}\cdot\nabla\psi\Big{|}\\ +\|D(u^{+}_{\tau_{n}})-D(u)\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{2}(\Omega))}\|\nabla c_{\tau_{n}}^{+}\cdot\nabla\psi\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{2}(\Omega))}.$ (4.15)
Since $\|D(u)\|_{\operatorname{L}_{\infty}(\Omega)}\leq C$ from (4.8e), (4.9c) and (4.6) we conclude (4.12a). Using similar techniques we can obtain (4.12b) and (4.12c). We omit the proof of (4.12d) as it is given in (Burman and Rappaz, 2003, Lemma 5.4, p1152). ∎

### v3

@@ 3.10. Lemma (energy estimates for the concentration)
$\frac{\tau}{2}\|d_{\tau}c^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\frac{1}{2}d_{\tau}\|c^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+D_{\min}|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\\ \leq\frac{1}{2}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\frac{1}{2}\|c^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+C|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}+\frac{D_{\min}}{2}|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}.$ (3.54)
After rearranging (3.54), multiplying the resulting inequality by $\tau$ and summing over k=1,2,,lk=1,2,...,l, for any $l\leq K$, we have
@@
[…]
@@
$\begin{split}\frac{1}{4}\max_{l=1,...,K}\|c^{l}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau\frac{D_{\min}}{2}\ \sum_{k=1}^{K}|c^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \\ \leq C\Big(1+\tau\sum_{k=1}^{K}\|d_{\tau}u^{k}\|^{2}_{\operatorname{L}_{2}(\Omega)}+\tau\sum_{k=1}^{K}|\phi^{k}|^{2}_{\operatorname{H}^{1}(\Omega)}\Big)\operatorname{e}\!^{\frac{T}{2}}\end{split}$ (3.56)
and from Lemmas 3.6 and 3.9 we conclude
@@
$(d_{\tau}c^{k},\chi)_{((\operatorname{H}^{1}(\Omega))^{\prime},\operatorname{H}^{1}(\Omega))}\leq\|D(u^{k})\nabla c^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\nabla\chi\|_{\operatorname{L}_{2}(\Omega)}\\ +\|d_{\tau}u^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\chi\|_{\operatorname{L}_{2}(\Omega)}+\|D_{1}(u^{k},c^{k-1})\nabla\phi^{k}\|_{\operatorname{L}_{2}(\Omega)}\|\nabla\chi\|_{\operatorname{L}_{2}(\Omega)}\\ \leq C(|c^{k}|_{\operatorname{H}^{1}(\Omega)}+|\phi^{k}|_{\operatorname{H}^{1}(\Omega)}+\|d_{\tau}u^{k}\|_{\operatorname{L}_{2}(\Omega)})\|\chi\|_{\operatorname{H}^{1}(\Omega)}$ (3.58)
@@
[…]
The required result follows after multiplying by $\tau$, summing over k=1,2,,Kk=1,2,...,K and noting (3.57) and Lemmas 3.6 and 3.9. ∎
Proof of Theorem 3.2. Theorem 3.2 is a direct consequence of Lemmas 3.5, 3.7 and 3.8. ∎
@@ 4. Weak convergence of the limits
$u^{-}_{\tau}(t)=u^{k-1},\quad u^{+}_{\tau}(t)=u^{k}\text{ for }t_{k-1}<t\leq t_{k}.$ (4.1)
We introduce the sequence $u^{k}$’s piecewise linear interpolant in time
@@
We similarly define $c^{\pm}_{\tau}(t),\hat{c}_{\tau}(t),\phi^{\pm}_{\tau}(t)$, and $\hat{\phi}_{\tau}(t)$. Next recalling (3.2), the terms on the left hand side of (3.1) can be reformulated as
$\begin{split}((m(c^{k-1}))^{-}{\textsl{H}}_{1}(u^{k},u^{k-1}),\chi)&+((m(c^{k-1}))^{+}{\textsl{H}}_{2}(u^{k},u^{k-1}),\chi)\\ &=30\left(m(c^{k-1})u^{k-1}u^{k}(u^{k}-1)^{2},\chi\right)\\ &\phantom{=}+30\left((m(c^{k-1}))^{+}(u^{k}-1)u^{k}(u^{k-1}-u^{k}),\chi\right).\end{split}$ (4.3)
@@
$(\sigma(u^{+}_{\tau}(t))\nabla\phi^{+}_{\tau}(t),\nabla\eta)=-(\hat{u}_{\tau}^{\prime}(t)+\sigma^{\prime}(u_{\tau}^{+}(t))\partial_{x_{1}}u_{\tau}^{+}(t),\eta),$ (4.4c)
@@
$\begin{split}({\mathcal{F}}(c_{\tau}^{-}(t),u_{\tau}^{-}(t),u_{\tau}^{+}(t)),\chi):=30\left(m(c_{\tau}^{-})u_{\tau}^{-}(t)u_{\tau}^{+}(t)(u_{\tau}^{+}(t)-1)^{2},\chi\right)\\ +30\left((m(c_{\tau}^{-}(t)))^{+}(u_{\tau}^{+}(t)-1)u_{\tau}^{+}(t)(u_{\tau}^{-}(t)-u_{\tau}^{+}(t)),\chi\right).\end{split}$ (4.5)
@@ 4.1. Lemma (stability bounds)
The first and second bounds in (4.6) follow from Lemmas 3.3 and 3.6. For the last bound in (4.6), we note from (4.2) that $\partial_{t}\hat{u}_{\tau}(t)=d_{\tau}u^{k}$ on $(t_{k-1},t_{k})$, k=1,2,,Kk=1,2,...,K and thus the result follows from Lemma 3.6.
The bounds in (4.7) and (4.8) follow using similar arguments and the results of Lemmas 3.10 and 3.9 respectively. ∎
@@ 4.2. Lemma (weak compactness)
There exist $(u,c,\phi)$ and $(\partial_{t}u,\partial_{t}c)$ such that for a sequence $(\tau_{n})_{n\in\mathbb{N}}$ of positive numbers with $\tau_{n}\to 0$ as $n\to\infty$, we have the following for $1\leq p<\infty$ if $\dim\Omega=2$ and $1\leq p<6$ if $\dim\Omega=3$, up to a subsequence, which for simplicity of presentation we again denote by $\tau_{n}$,
@@ 4.3. Lemma (strong convergence)
For $u\in\operatorname{L}_{\infty}([0,T];\operatorname{L}_{\infty}(\Omega))$ and $c\in\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega))$, with $1\leq p<\infty$ if $\dim\Omega=2$ and $1\leq p<6$ if $\dim\Omega=3$, we have the following as $\tau_{n}\to 0$,
@@
$D_{1}(u^{+}_{\tau_{n}},c^{-}_{\tau_{n}})\to D_{1}(u,c)\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega)),$ (4.10f)
$30\,m(c_{\tau}^{-})u_{\tau}^{-}u_{\tau}^{+}(u_{\tau}^{+}-1)^{2}\to m(c)h^{\prime}(u)\text{ in }\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega)),$ (4.10g)
@@
Noting that $g^{\prime}(u),D(u),\sigma(u)$ are polynomials in $u$ and $m(c)$ is linear in $c$, the proofs of (4.10a)–(4.10e) follow from Lemma 4.1, (4.9a), (4.9d), (4.9h) and (4.9i). Noting, from definition (2.10), that $D_{1}$ is a bounded Lipschitz function we obtain
$\|D_{1}(u_{\tau}^{+},c_{\tau}^{-})-D_{1}(u,c)\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega))}\\ \leq C(\|u_{\tau}^{+}-u\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega))}+\|c_{\tau}^{-}-c\|_{\operatorname{L}_{2}([0,T];\operatorname{L}_{p}(\Omega))})$ (4.11)
and (4.10f) then follows by using (4.9h) and (4.9i). Adopting similar techniques and noting from (2.12) that $h^{\prime}(u)=30u^{2}(u-1)^{2}$, from (3.14), (4.6), (4.10b) and (4.10e) we conclude (4.10g).
@@ 4.4. Lemma (weak convergence)
For some $1\leq r\leq 2p/(p+2)$, with $1\leq p<\infty$ if $\dim\Omega=2$ and $1\leq p<6$ if $\dim\Omega=3$, we have the following as $\tau_{n}\to 0$
$D(u^{+}_{\tau_{n}})\nabla c^{+}_{\tau_{n}}\rightharpoonup D(u)\nabla c\text{ in }\operatorname{L}_{1}([0,T];(\operatorname{L}_{r}(\Omega))^{d}),$ (4.13a)
$D_{1}(u^{+}_{\tau_{n}},c^{-}_{\tau_{n}})\nabla\phi^{+}_{\tau_{n}}\rightharpoonup D_{1}(u,c)\nabla\phi\text{ in }\operatorname{L}_{1}([0,T];(\operatorname{L}_{r}(\Omega))^{d}),$ (4.13b)
$\sigma(u^{+}_{\tau_{n}})\nabla\phi_{\tau_{n}}^{+}\rightharpoonup\sigma(u)\nabla\phi\text{ in }\operatorname{L}_{1}([0,T];(\operatorname{L}_{r}(\Omega))^{d}),$ (4.13c)
@@
From Lemmas 4.3 and 4.2, for $1\leq p<\infty$ if $\dim\Omega=2$ and $1\leq p<6$ if $\dim\Omega=3$, we have
$\begin{split}D(u_{\tau_{n}}^{+})\to D(u)\text{ in }\operatorname{L}_{2}(0,T;\operatorname{L}_{p}(\varOmega))\\ \nabla c_{\tau_{n}}^{+}\rightharpoonup\nabla c\text{ in }\operatorname{L}_{2}(0,T;(\operatorname{L}_{2}(\varOmega))^{d}).\end{split}$ (4.14)
From the extended result of Boyer and Fabrie (2013, Prop. II.2.30 on p64 and p92–93) we have that
$D(u_{\tau_{n}}^{+})\nabla c_{\tau_{n}}^{+}\rightharpoonup D(u)\nabla c\text{ in }\operatorname{L}_{1}(0,T;(\operatorname{L}_{r}(\varOmega))^{d})\text{ with }r=\frac{2p}{p+2},$ (4.15)
which establishes (4.13a). Similar arguments lead to (4.13c) and (4.13b). Relation (4.13d) follows immediately from (Burman and Rappaz, 2003, Lemma 5.4, p1152). ∎